# Learning Noise-Robust Stable Koopman Operator for Control with Physics-Informed Observables

### v2

@@ II-E Curriculum Learning
Curriculum learning [43] is a training strategy where a model is first trained on simpler data points and then gradually introduced to more challenging ones. This approach helps the model to progressively learn and adapt to increasing complexity.
@@
In our framework, we implement a form of curriculum learning by progressively increasing the roll-out length $R$ during the training process. Initially, the model is trained with shorter roll-out lengths, focusing on immediate predictions. As the training progresses and the model stabilizes, the roll-out length $R$ is increased geometrically after certain epochs. This gradual increase is expected to allow the model to learn from easier short-term predictions before tackling the more complex task of long-term prediction, thereby aligning with the curriculum learning paradigm. By doing so, the model is expected to build its capacity to handle extended sequences, improving overall prediction accuracy over long time horizons.
II-F Stabilized Formulation
To promote stability of the lifted system, we enforce stable eigenvalues of $A$ by employing the fact that a skew symmetric matrix with negative real valued diagonal elements have stable or neutral stable eigenvalues in continuous time dynamical system [31]. For any $Q\in\mathbb{R}^{\tilde{n}\times\tilde{n}}$ matrix and $C\in\mathbb{R}^{\tilde{n}\times\tilde{n}}$ diagonal matrix with negative real diagonal elements, $\tilde{A}=Q-Q^{T}+C$ has eigenvalues in the negative real part of the complex plane. We compute the discrete-time $\hat{A}\in\mathbb{R}^{\tilde{n}\times\tilde{n}}$ by taking the matrix exponentiation of $\tilde{A}=Q-Q^{T}+C$ multiplied with the sampling timestep $\Delta t$.
$\hat{A}$ $=\exp\left(\tilde{A}\Delta t\right)$ (14)
@@ II-F Stabilized Formulation
Note that a similarity transform of $\hat{A}$ via the coordinate transform matrix $P\in\mathbb{R}^{\tilde{n}\times\tilde{n}}$ is required to recover $A$.
$A$ $=P\hat{A}P^{-1}$ (15)
@@
This formulation ensures the stability of the Koopman Operator throughout the training, which we refer as Dissipative Koopman Operator. In contrast, we refer Standard Koopman Operator as lifted system with an unconstrained matrix $A$.
II-G Model Predictive Control
@@
$\min_{u_{k},z_{k}}$ $J\left(\{z_{k}\}_{k=0}^{N_{p}},\{u_{k}\}_{k=0}^{N_{p}}\right),$ (16)
subject to $\quad z_{k+1}=Az_{k}+Bu_{k},\quad k=0,1,\ldots,N_{p}-1,$
$\quad e_{k}=r_{k}-Cz_{k},$
@@
where $u_{\min/\max}$ are the input bounds, $z_{\min/\max}$ are the bounds on the lifted states, and $r_{k}$ is the reference signal. Only the first control input $u_{0}$ from the optimal sequence is applied, and the optimization problem is resolved at the next time step. The cost function $J$ is given by:
$J(z,u)=e_{N_{p}}^{\top}Q_{N}e_{N_{p}}+\sum_{k=0}^{N_{p}-1}\left(e_{k}^{\top}Qe_{k}+u_{k}^{\top}Ru_{k}\right).$ (17)
II-H Iterative Learning
Models trained on trajectories that are generated by randomly sampling the phase space and control input could perform poorly when subjected to the optimal feedback control inputs. This is primarily because these trained models lack exposure to the desired scenarios during the training process, especially for relatively high-dimensional system. To address this issue, we propose a straightforward iterative data augmentation technique inspired by the work of Uchida and Duraisamy [44]. The idea is to iteratively collect response of the true system under control input synthesized with the previous data-driven model. Such response together with control input is used to augment the original dataset to further refine and improve the data-driven model. Our rationale is that such experience from the true system can be used to better explore the phase space of nonlinear system than merely ad-hoc random sampling the phase space, which is subject to the curse of dimensionality. Hence, our proposed framework is named as STIRK (Stable Training with Iterative Roll-out Koopman Operator). A summary of the proposed framework is illustrated in the figure 1.
@@ III Numerical Experiments
In this section, we will evaluate our proposed framework against other popular frameworks by considering the following two tasks:
  • Prediction of the time evolution of Van der Pol oscillator given initial condition using the surrogate model learned with our framework;
  • Model predictive control of CartPole system given a surrogate model learned with our framework from noisy measurements.
III-A Prediction Performance
@@
Van der Pol Oscillator.
$\dot{x}_{1}$ $=-x_{2},$ (18)
$\dot{x}_{2}$ $=\mu\left(-1+x_{1}^{2}\right)x_{2}+x_{1},$ (19)
where $\mu$ is set to 1.0. The Van der Pol Oscillator is a nonlinear system characterized by a non-conservative force that can either dampen or amplify oscillations depending on the state. Since our framework relies on discrete-time dynamical system, the continuous-time version of Van der Pol system is discretized using a fourth-order Runge-Kutta method with a time step $\Delta t$ of 0.1. The training dataset is generated by running simulations for 100 steps, also with a time step of $\Delta t$ of 0.1. After the simulation, Gaussian noise is added to the states, with 10 noise levels logarithmically spaced from $10^{-3}$ to $10^{-1}$. For models trained on a single trajectory, the initial condition is sampled from a normal distribution with a mean of $[-1.0,-1.0]$ and a standard deviation of $[0.05,0.05]$. For models trained on multiple trajectories, 50 initial conditions are uniformly sampled from the range $[-1,1]$. To ensure robustness and generalizability, we create 10 sets of training data using different random seeds. Figure 2 illustrates one set of the training trajectories and test trajectories within the $x_{1},x_{2}$ plane.
Prediction Performance Evaluation
To evaluate the prediction performance of the different algorithms, we use a normalized error metric that quantifies the discrepancy between the predicted trajectories and the ground truth trajectories. Let $\mathbf{x}(t)\in\mathbb{R}^{n}$ represent the true state of the system at time $t$ and $\hat{\mathbf{x}}(t)\in\mathbb{R}^{n}$ represent the predicted state at time $t$. The normalized prediction error over the trajectory is computed using the following formula:
@@
$\mathbf{X}=\begin{bmatrix}\mathbf{x}(1)\\ \mathbf{x}(2)\\ \vdots\\ \mathbf{x}(T)\end{bmatrix},\quad\hat{\mathbf{X}}=\begin{bmatrix}\hat{\mathbf{x}}(1)\\ \hat{\mathbf{x}}(2)\\ \vdots\\ \hat{\mathbf{x}}(T)\end{bmatrix}.$ (21)
@@
The Van der Pol system was initially trained using the Adam optimizer with a base learning rate of 0.01. A cyclic learning rate scheduler in triangular mode was employed, where the initial learning rate was halved each cycle. The maximum learning rate was set to 0.1, with each cycle stepping up for 500 epochs and stepping down for another 500 epochs. The Polyflow order was configured to 4. The training process spanned 15,000 epochs with a batch size of 1,000. The roll-out length was progressively increased, starting at 1 and doubling every 200 epochs, with a maximum roll-out length initially set to 90. In the final 1,000 epochs, the optimizer was switched to LBFGS with a learning rate of 0.01, applied on full batch data, to fine-tune the system.
@@
Figure 3a compares the reconstruction performance of various Dynamic Mode Decomposition (DMD)-based methods when trained on a single trajectory of the Van der Pol system at a noise level of 0.0599. The legend lists the normalized error for each method. The Dissipative Koopman model with Polyflow basis functions (DISSIPATIVE-Polyflow) achieved the lowest error of $2.98\times 10^{-2}$, indicating superior accuracy. OptDMD, using both identity and Polyflow basis functions, performed similarly well, closely approaching the performance of the Dissipative Koopman model. However, since OptDMD also optimizes the initial condition, it shows a deviation from the actual initial condition. Other methods exhibited higher errors, indicating less accurate reconstructions under the given noise conditions. Figure 3b presents the mean squared error (MSE) trajectory over 100 timesteps for the Van der Pol system at a noise level of 0.0599 for 10 different training sets. Each line represents the mean of the errors for a different DMD-based method, with shaded areas indicating the standard deviation of the error. The Dissipative Koopman model with Polyflow basis functions consistently maintains the lowest MSE across all timesteps, highlighting its robustness and accuracy under noisy conditions. The eDMD with Polyflow basis functions and the eDMD with RBF basis functions exhibit higher MSEs, particularly in the early timesteps, indicating less accurate predictions compared to the Dissipative model. The standard DMD with identity observables (DMD-Identity) also shows higher error compared to the Dissipative Koopman model but performs better than some of the other methods. The figures demonstrate that the Dissipative Koopman model with Polyflow basis functions outperforms other DMD-based methods in terms of prediction accuracy for the Van der Pol system under noisy conditions. This highlights the effectiveness of incorporating dissipative properties and structured observables in enhancing the robustness and accuracy of the model.
@@
III-B Control Performance
CartPole System
@@ III-B Control Performance
The CartPole system consists of a cart that moves along a track while balancing a pole attached to it via a pivot. The dynamics of the system are described by the equations above, where $x_{1}$ represents the position of the cart, $x_{2}$ represents the angle of the pole, $x_{3}$ represents the velocity of the cart, $x_{4}$ represents the angular velocity of the pole, and $u$ is the control input force applied to the cart. The parameters of the system include the mass of the cart ($m_{c}$), the mass of the pole ($m_{p}$), the length of the pole ($l$), and the acceleration due to gravity ($g$). For our framework, the continuous-time CartPole system is discretized using a fourth-order Runge-Kutta method with a time step $\Delta t$ of 0.05. The training dataset is generated by running simulations for 200 steps and contains 50 different trajectories. After the simulation, Gaussian noise is added to the states with noise level of 0.1. The initial conditions for the CartPole system are randomly sampled with the position ($x_{1}$) from $[-1,1]$, the angle ($x_{2}$) from $[-\pi/2,\pi/2]$, the velocity ($x_{3}$) from $[-0.1,0.1]$, and the angular velocity ($x_{4}$) from $[-0.1,0.1]$. Control inputs are generated using exponentially decaying sine waves of randomly sampled frequencies, amplitudes and phase angles.
@@ IV Iterative Data Augmentation
To more effectively explore the phase space of a nonlinear system, we introduce an iterative data augmentation technique. Instead of relying solely on ad-hoc random sampling, we augment the training data with failure trajectories from closed-loop control and then retrain the Koopman operator. This approach addresses challenges highlighted by [49], where the authors emphasize the need to identify inputs that closely approximate the final closed-loop input and to use randomized signals that resemble it. However, finding suitable randomized control signals can be challenging and may not be adequate for control tasks. Therefore, we propose a rather automated approach: augmenting the training data with control signals from previous failed control tasks in a closed-loop setting. This approach involves collecting the system’s response to control inputs from data-driven based model predictive controller. The responses, along with the corresponding control inputs, are used to enhance the original dataset, enabling further refinement and improvement of the model.
@@
Figure 6 visualizes the trajectory of the CartPole system across three different scenarios: no data augmentation, one-time data augmentation, and twice data augmentation. In the first scenario, where no data augmentation was applied, the Koopman model fails to stabilize the CartPole after 200 steps. However, when the failed trajectory, along with other failed trajectories from different initial conditions, is added to the training data, the Koopman model demonstrates significantly improved control outcomes. The performance notably improves with data augmentation, transitioning from oscillating to stable trajectories. This enhancement is also reflected in the reduction of the Model Predictive Control (MPC) cost, with costs decreasing from 135.6 (no augmentation) to 74.7 (one-time augmentation) and 67.8 (twice augmentation).
@@
Figure 7 presents violin plots illustrating the distribution of Model Predictive Control (MPC) costs for the CartPole system across two different tasks—seen and unseen—and iterative models under a noise level of 0.1. The “seen” task refers to scenarios where input bounds are set between $[-20,20]$, while the “unseen” task refers to tighter input bounds of $[-10,10]$. Data augmentation was performed using closed loop trajectories from the seen task. It is important to note that classical iterative learning [50] cannot easily generalize to a different task. The violin plots display MPC costs for four different initial states and two levels of data augmentation, comparing the costs for both control tasks. Each subplot emphasizes the performance gains achieved through iterative learning. The results indicate that both seen and unseen control tasks benefit significantly from iterative learning, with successive iterations leading to lower and more stable MPC costs. This demonstrates the effectiveness of iterative learning in enhancing the overall performance of the system.

### v5

@@ II-F Curriculum Learning
Curriculum learning [45] is a training strategy in which a model is first trained in simpler tasks and then gradually exposed to more challenging ones. This approach helps the model to learn and adapt progressively to increasing complexity.
@@
In our framework, we implement a form of curriculum learning by progressively increasing the length of the rollout $R$ during the training process. Initially, the model is trained with shorter rollout lengths, focusing on immediate predictions. As training progresses and the model stabilizes, the roll-out length $R$ increases geometrically after certain epochs. This gradual increase is expected to allow the model to learn from easier short-term predictions before tackling the more complex task of long-term prediction, thus aligning with the curriculum learning paradigm. In doing so, the model is expected to build its capacity to handle extended sequences, improving overall prediction accuracy over long time horizons. Compared to direct exposure to long rollout lengths from the start, this progressive approach offers optimization benefits by improving stability and convergence. Training with long roll-outs early on can lead to high variance in gradients and poor local optima due to compounding prediction errors. By starting with shorter roll-outs, the model establishes a more robust foundation before incrementally handling extended sequences, ultimately enhancing prediction accuracy over long time horizons.
II-G Stabilized Formulation
To guarantee the stability of both the Koopman matrix $A$ and the training process under rollout loss, we enforce stable eigenvalues of $A$ employing the fact that a skew-symmetric matrix with negative real-valued diagonal elements has stable or neutral stable eigenvalues in continuous-time dynamical system [34]. This guarantees the stability of the Koopman matrix during training with roll-out loss, preventing instability in both forward and backward passes.
For some matrix $Q\in\mathbb{R}^{\tilde{n}\times\tilde{n}}$, $\sigma_{1},...,\sigma_{\tilde{n}}\in\mathbb{R}$, a diagonal matrix $C=\mathrm{diag}(-\sigma_{1}^{2},...,-\sigma_{\tilde{n}}^{2})\in\mathbb{R}^{\tilde{n}\times\tilde{n}}$ with negative real diagonal elements, we construct:
$\tilde{A}=Q-Q^{\top}+C.$ (13)
By construction, the eigenvalues of $\tilde{A}$ lie in the left half of the complex plane, ensuring stability for the continuous-time dynamical system.
To obtain the discrete-time matrix $\hat{A}\in\mathbb{R}^{\tilde{n}\times\tilde{n}}$, we apply the matrix exponential to $\tilde{A}$ scaled by the sampling timestep $\Delta t$:
$\hat{A}=\exp\left(\tilde{A}\Delta t\right).$ (14)
@@ II-G Stabilized Formulation
Since the Koopman operator is typically expressed in a different coordinate system, a similarity transformation using a coordinate transformation matrix $P\in\mathbb{R}^{\tilde{n}\times\tilde{n}}$ is required to recover $A$:
$A=P\hat{A}P^{-1}.$ (15)
@@
This formulation ensures that the Koopman operator remains stable throughout the training, which we refer to as the dissipative Koopman operator. In contrast, the standard Koopman operator refers to a lifted system where $A$ is unconstrained.
II-H Model Predictive Control
@@
$\min_{\{u_{k}\}_{k=0}^{N_{p}-1}}$ $e_{N_{p}}^{\top}Q_{N}e_{N_{p}}+\sum_{k=0}^{N_{p}-1}\left(e_{k}^{\top}Qe_{k}+u_{k}^{\top}Ru_{k}\right),$ (16)
subject to $\quad\quad e_{k}=r_{k}-Cz_{k},$
$z_{k+1}=Az_{k}+Bu_{k},\quad k=0,1,\ldots,N_{p}-1,$
@@
where $u_{\min/\max}$ are the input bounds, $z_{\min/\max}$ are the bounds on the lifted states, and $r_{k}$ is the reference signal. Only the first control input $u_{0}$ from the optimal sequence is applied, and the optimization problem is resolved at the next time step.
II-I Iterative Learning
Models trained on trajectories that are generated by randomly sampling the phase space and control input could perform poorly when subjected to optimal feedback control input. This is mainly because these trained models lack exposure to the desired scenarios during the training process, especially for a relatively high-dimensional system. To address this issue, we propose a straightforward iterative data augmentation technique inspired by the work of Uchida and Duraisamy [46]. The idea is to iteratively collect the response of the true system under control input synthesized with the previous data-driven model. Such response together with control input is used to augment the original dataset to further refine and improve the data-driven model. Our rationale is that such experience from the true system can be used to better explore the phase space of the nonlinear system than merely ad hoc random sampling of the phase space, which is subject to the curse of dimensionality. Hence, our proposed framework is named as STIRK (Stable Training with Iterative Roll-out Koopman Operator). A summary of the proposed framework is illustrated in the figure 1.
@@ III Numerical Experiments
In this section, we evaluate our proposed framework against other popular approaches in five different systems.
  1. 1.
    Autonomous Van der Pol Oscillator – Used to compare prediction performance.
  2. 2.
    Cartpole System – Evaluated on the task of stabilizing the pole in the upright position.
  3. 3.
    Mountain Car (continuous action space version) from [47]
  4. 4.
    Lunar Lander (continuous action space version) from [47]
  5. 5.
    Panda Emika Robot from [48]
While the first two examples assume known system dynamics, the last three environments are evaluated under the assumption of unknown system dynamics.
III-A Autonomous Van der Pol Oscillator
$\dot{x}_{1}$ $=-x_{2},$ (17)
$\dot{x}_{2}$ $=\mu\left(-1+x_{1}^{2}\right)x_{2}+x_{1},$ (18)
where $\mu$ is set to 1.0. The Van der Pol Oscillator is a nonlinear system characterized by a non-conservative force that can either dampen or amplify oscillations depending on the state. Since our framework relies on the discrete-time dynamical system, the continuous-time version of the van der Pol system is discretized using a fourth-order Runge-Kutta method with a time step $\Delta t$ of 0.1. The training dataset is generated by running simulations for 100 steps, also with a time step of $\Delta t$ of 0.1. After simulation, Gaussian noise is added to the states, with 10 logarithmically spaced noise levels from $10^{-3}$ to $10^{-1}$. For models trained on a single trajectory, the initial condition is sampled from a normal distribution with a mean of $[-1.0,-1.0]$ and a standard deviation of $[0.05,0.05]$. For models trained on multiple trajectories, 50 initial conditions are uniformly sampled from the range $[-1,1]$. To ensure robustness and generalizability, we created 10 sets of training data using different random seeds. Figure 2 illustrates one set of training trajectories and test trajectories within the $x_{1},x_{2}$ plane.
Prediction Performance Evaluation
@@ III-A Autonomous Van der Pol Oscillator
To evaluate the prediction performance of these different algorithms, we use a normalized error metric that quantifies the discrepancy between the predicted trajectories and the ground truth trajectories. Let $\mathbf{x}(t)\in\mathbb{R}^{n}$ represent the true state of the system at time $t$ and $\hat{\mathbf{x}}(t)\in\mathbb{R}^{n}$ represent the predicted state at time $t$. The normalized prediction error over the trajectory is computed using the following formula:
@@
$\mathbf{X}=\begin{bmatrix}\mathbf{x_{1}}\\ \mathbf{x_{2}}\\ \vdots\\ \mathbf{x_{T}}\end{bmatrix},\quad\hat{\mathbf{X}}=\begin{bmatrix}\hat{\mathbf{x}}_{1}\\ \hat{\mathbf{x}}_{2}\\ \vdots\\ \hat{\mathbf{x}}_{T}\end{bmatrix}.$ (20)
@@
The Van der Pol system was initially trained using the Adam optimizer with a base learning rate of 0.01. A triangular-mode cyclic learning rate scheduler was employed, where the initial learning rate was halved each cycle. The maximum learning rate was set to 0.1, and each cycle increased for 500 epochs and decreased for another 500 epochs. The Polyflow order was configured to 4. The training process spanned 15,000 epochs with a batch size of 1,000. The rollout length was progressively increased, starting at 1 and doubling every 200 epochs, with a maximum rollout length initially set at 90. In the final 1,000 epochs, the optimizer was switched to L-BFGS with a learning rate of 0.01, applied on full batch data, to fine-tune the system.
@@
Figure 3a compares the reconstruction performance of various methods based on dynamic mode decomposition (DMD) when trained on a single trajectory of the Van der Pol system at a noise level of 0.0599. The legend lists the normalized error for each method. The dissipative Koopman model with polyflow basis functions (Dissipative-Polyflow) achieved the lowest error of $2.98\times 10^{-2}$, indicating a superior precision. OptDMD, using both identity and Polyflow basis functions, performed similarly well, closely approaching the performance of the dissipative Koopman model. However, since OptDMD also optimizes the initial condition, it shows a deviation from the actual initial condition. Other methods exhibited higher errors, indicating less accurate reconstructions under the given noise conditions. Figure 3b presents the mean squared error (MSE) trajectory over 100 time steps for the Van der Pol system at a noise level of 0.0599 for 10 different training sets. Each line represents the mean error for a different DMD-based method, with shaded areas indicating the standard deviation of the error. The dissipative Koopman model with polyflow basis functions consistently maintains the lowest MSE across all timesteps, highlighting its robustness and accuracy under noisy conditions. The eDMD with polyflow basis functions and the eDMD with RBF basis functions exhibit higher MSEs, particularly in the early timesteps, indicating less accurate predictions compared to the dissipative model. The standard DMD with identity observables (DMD-Identity) also shows higher error compared to the dissipative Koopman model but performs better than some of the other methods. The figures demonstrate that the dissipative Koopman model with Polyflow basis functions outperforms other DMD-based methods in terms of prediction accuracy for the Van der Pol system under noisy conditions. This highlights the effectiveness of incorporating dissipative properties and structured observables in enhancing the robustness and accuracy of the model.
@@
III-B CartPole System
@@
The CartPole system consists of a cart that moves along a track, balancing a pole attached to it via a pivot. The dynamics of the system are described by the equations above, where $x_{1}$ represents the position of the cart, $x_{2}$ represents the angle of the pole, $x_{3}$ represents the velocity of the cart, $x_{4}$ represents the angular velocity of the pole, and $u$ is the control input force applied to the cart. The parameters of the system include the mass of the cart ($m_{c}$), the mass of the pole ($m_{p}$), the length of the pole ($l$), and the acceleration due to gravity ($g$). For our framework, the continuous-time CartPole system is discretized using a fourth-order Runge-Kutta method with a time step $\Delta t$ of 0.05. The training dataset is generated by running simulations for 200 steps and contains 50 different trajectories. After the simulation, Gaussian noise is added to the states with noise level of 0.1. The initial conditions for the CartPole system are randomly sampled with the position ($x_{1}$) from $[-1,1]$, the angle ($x_{2}$) from $[-\pi/2,\pi/2]$, the velocity ($x_{3}$) from $[-0.1,0.1]$, and the angular velocity ($x_{4}$) from $[-0.1,0.1]$. Control inputs are generated using exponentially decaying sine waves of randomly sampled frequencies, amplitudes and phase angles.
@@ Results
III-C Mountain Car
Mountain Car Continuous Version 0 is an environment from the Gymnasium library [47] where a car is tasked with reaching the top of a mountain. The control action is a force applied along the x-axis with bounds $[-1.0,2.0]$. The observed state consists of the car’s position along the x-axis and its velocity.
We apply our learning framework to this environment under the assumption that the governing equation is unknown, and we have access only to noisy measurements of state-action sequences. For data generation, we use uniformly distributed random actions within the action space bounds and uniformly distributed initial positions between $[-0.6,-0.4]$ with zero initial velocity. The episode terminates upon reaching the target or after 1000 steps, with each step representing 0.02s. We collect 200 episodes of varying step lengths.
To preprocess the dataset, we normalize the states and actions by subtracting their mean values and dividing by their standard deviations. Additionally, we introduce Gaussian noise with mean 00 and variances $\sigma^{2}=\{0.01,0.1,0.5\}$ to create three different training sets. A separate test dataset is generated following the same procedure.
(a) Mountain Car Test Prediction Error
(b) Lunar Lander Test Prediction Error
(c) Panda Test Prediction Error
Figure 6: Test trajectory prediction errors for the Mountain Car, Lunar Lander, and Panda environments under unknown dynamics. For these experiments, a Polyflow order of 2 was used for the lifted state construction, and all models were trained with training data subject to 0.1 noise levels. Optimization was performed using the SOAP optimizer [53].
III-D Lunar Lander
The Lunar Lander environment is a Gymnasium environment [47] in which an agent controls a lander’s thrusters to achieve a soft landing on a designated pad. The state space originally comprises the lander’s position, velocity, angle, and angular velocity, while the action space consists of continuous thrust controls.
(a) Mountain Car MPC Trajectory
(b) Lunar Lander MPC Trajectory
(c) Panda MPC Trajectory
Figure 7: MPC-controlled trajectories for Mountain Car, Lunar Lander and Panda environments, trained on data with a noise level of 0.1. Model predictive control was implemented using the do-mpc [54] package.
III-E Panda Emika Franka Robot
The Panda Emika Franka robotic arm, as provided in the panda-gym environment [48], is a 7-DoF manipulator widely used in industrial and research applications. In our experiments, the objective is to achieve precise target positioning by applying continuous control actions via a model predictive control (MPC) framework. The state space consists of the position and velocity of the robot’s end effector, while the control inputs correspond to the forces applied at the end effector.
Figure 8: Effectiveness of dissipative roll-out under noise. Heatmaps show predictive test performance across different noise levels and sequence lengths.
@@ IV Iterative Data Augmentation
To explore the phase space of a nonlinear system more effectively, we introduce an iterative data augmentation technique. Instead of relying solely on ad hoc random sampling, we augment the training data with failure trajectories from closed-loop control and then retrain the Koopman operator. This approach addresses challenges highlighted by [55], where the authors emphasize the need to identify inputs that closely approximate the final closed-loop input and to use randomized signals that resemble it. However, finding suitable randomized control signals can be challenging and may not be adequate for control tasks. Therefore, we propose a systematic approach: augmenting the training data with control signals from previous failed control tasks in a closed-loop setting. This approach involves collecting the system’s response to control inputs from data-driven based model predictive controller. The responses, along with the corresponding control inputs, are used to enhance the original data set, allowing further refinement and improvement of the model.
@@
Figure 9 visualizes the trajectory of the CartPole system in three different scenarios: no data augmentation, one-time data augmentation, and twice data augmentation. In the first scenario, where no data augmentation was applied, the Koopman model fails to stabilize the CartPole after 200 steps. However, when the failed trajectory, along with other failed trajectories from different initial conditions, is added to the training data, the Koopman model demonstrates significantly improved control outcomes. Performance notably improves with data augmentation, transitioning from oscillating to stable trajectories. This enhancement is also reflected in the reduction of the Model Predictive Control (MPC) cost, with costs decreasing from 135.6 (no augmentation) to 74.7 (one-time augmentation) and 67.8 (twice augmentation).
@@
Figure 10 presents violin plots that illustrate the distribution of Model Predictive Control (MPC) costs for the CartPole system on two different tasks, seen and unseen, and iterative models under a noise level of 0.1. The “seen” task refers to scenarios where input bounds are set between $[-20,20]$, while the “unseen” task refers to tighter input bounds of $[-10,10]$. Data augmentation was performed using closed-loop trajectories from the seen task. It is important to note that iterative classical learning [56] cannot be easily generalized to a different task. The violin plots display MPC costs for four different initial states and two levels of data augmentation, comparing the costs for both control tasks. Each subplot emphasizes the performance gains achieved through iterative learning. The results indicate that both seen and unseen control tasks benefit significantly from iterative learning, with successive iterations leading to lower and more stable MPC costs. This shows the effectiveness of iterative learning in improving the overall performance of the system.